# Entanglement, defects, and $T\bar{T}$ on a black hole background

### v1

@@ 2.4 $T\bar{T}$ deformation
We now briefly review $T\bar{T}$ deformation in CFTs and its holographic dual. The $T\bar{T}$ deformation of a $2d$ conformal field theory considered on a flat space is defined as [38, 39]
$\frac{dI_{QFT}^{(\mu)}}{d\mu}=-2\pi\int d^{2}x(T\bar{T})_{\mu},\qquad I_{QFT}^{(\mu)}\Bigg|_{\mu=0}=I_{CFT},$ (2.15)
@@
$(T\bar{T})=\frac{1}{8}\left(T_{ab}T^{ab}-(T^{a}_{a})^{2}\right).$ (2.16)
This operator is shown to satisfy the factorization formula as [37]
$\langle T\bar{T}\rangle=\frac{1}{8}\left(\langle T_{ab}\rangle\langle T^{ab}\rangle-\langle T_{a}^{a}\rangle^{2}\right).$ (2.17)
The corresponding bulk dual of the $T\bar{T}$ deformed CFT2 in a flat space was proposed by the authors in [70, 50] to be a quantum gravity in AdS3 with a radial cut-off as
$ds^{2}=\frac{1}{z^{2}}\left(-dt^{2}+dx^{2}+dz^{2}\right),\qquad z>z_{c},$ (2.18)
where Dirichlet boundary conditions are imposed at $z=z_{c}$. The deformation parameter is related to this radial cut-off $z_{c}$ as[50]
$\mu=8G_{N}z_{c}^{2}=\frac{12}{c}z_{c}^{2}.$ (2.19)
@@ 3.1 Including brane defects
$ds^{2}=d\rho^{2}+\frac{\cosh^{2}\rho}{u^{2}}\bigg[\left(1-\frac{u}{u_{h}}\right)d\tau_{bs}^{2}+\frac{du^{2}}{\left(1-\frac{u}{u_{h}}\right)}\bigg],$ (3.1)
@@
$\omega=\frac{u_{h}e^{\frac{i\tau_{bs}}{2u_{h}}}}{\sqrt{1-\frac{u}{u_{h}}}},$ (3.2)
@@
$ds^{2}=d\rho^{2}+\frac{4u_{h}^{2}\cosh^{2}\rho}{(u_{h}^{2}-\omega\bar{\omega})^{2}}d\omega d\bar{\omega}.$ (3.3)
@@
$\omega=\frac{u_{h}}{\tau_{p}+iy-\frac{i}{2}}-iu_{h},$ (3.4)
@@
$ds^{2}=d\rho^{2}+\Omega_{y}^{2}(d\tau_{p}^{2}+dy^{2}),\qquad\Omega_{y}=\Big|\frac{\cosh\rho}{y}\Big|,$ (3.5)
which is the same as eq. 2.3. The conformal boundary is now located at $y=0$. Finally using the transformations
[…]
@@ 3.2 Entanglement Entropy
We now compute the entanglement entropy for a subsystem $A\equiv[L,u_{h}]\cup[u_{h},R]$ on a constant time slice in the radiation baths of the eternal black hole model described previously. The endpoints of $A$ are denoted by the points $P\equiv(t,L,\rho_{\epsilon})$ and $Q\equiv(-t+i2\pi u_{h},R,\rho_{\epsilon})$ on the asymptotic boundary. [^4] Note that to account for the reversal of the timelike Killing vector as one crosses black string horizon from the left to the right partition, we take the time coordinate on the right partition $t_{R}$ to be related to that of the left partition $t_{L}$ as $t_{R}\to-t_{L}+i2\pi u_{h}$. For simplicity of computation we begin in the Poincaré coordinates, where the corresponding endpoints are now at $P\equiv(\tau_{1},x_{1},z_{\epsilon_{1}})$ and $Q\equiv(\tau_{2},x_{2},z_{\epsilon_{2}})$ such that $z_{\epsilon_{1}},z_{\epsilon_{2}}\to 0$. We later transform the final result using eqs. 3.2, 3.4 and 3.6 to express the entanglement entropy in the AdS black string coordinates once again. Depending on the distance of subsystem $A$ from the EOW brane considered at a constant $\rho=-\rho_{b}$, we come across two distinct entanglement entropy phases, which are discussed below.
@@ 3.2.1 Island Phase
Field theory computation : In this phase the points $P$ and $Q$ are assumed to be close to the brane such that the subsystem $A$ on the radiation bath has a corresponding island region on the brane, as illustrated in fig. 2. Under these circumstances the entanglement entropy may be determined using the island formula in eq. 2.8, where the effective entanglement entropy may be expressed in terms of the 4-point correlators of twist fields at the points $P,P^{\prime},Q$ and $Q^{\prime}$, which in the large $c$ limit factorizes as
$\langle\sigma_{n}(P)\bar{\sigma}_{n}(P^{\prime})\sigma_{n}(Q^{\prime})\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}\approx\langle\sigma_{n}(P)\bar{\sigma}_{n}(P^{\prime})\rangle_{{}_{\text{CFT}}}\langle\sigma_{n}(Q^{\prime})\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}.$ (3.8)
This allows us to express the first term in eq. 2.8 as
$S_{eff}=\lim_{n\to 1}\frac{1}{1-n}\Big[\log\Omega_{P^{\prime}}^{-2h_{n}}\langle\sigma_{n}(P)\bar{\sigma}_{n}(P^{\prime})\rangle_{{}_{\text{CFT}}}+\log\Omega_{Q^{\prime}}^{-2h_{n}}\langle\sigma_{n}(Q^{\prime})\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}\Big],$ (3.9)
where $\Omega_{P^{\prime}}$ and $\Omega_{Q^{\prime}}$ are the conformal factor at the points $P^{\prime}$ and $Q^{\prime}$ respectively, which are determined using eq. 3.5. Assuming $P^{\prime}\equiv(-y_{{}_{P^{\prime}}},\tau_{{}_{P^{\prime}}})$ and $Q^{\prime}\equiv(-y_{{}_{Q^{\prime}}},\tau_{{}_{Q^{\prime}}})$, with the coordinate $y$ running along the brane, the effective entanglement entropy may be obtained using the standard form of the 2-point correlator as
@@
[…]
where $\epsilon_{y}$ denotes the UV cut-off on the brane. Adding to it the area term provided in [22], the generalized entanglement entropy in this scenario may then be expressed as
$S_{gen}=S_{eff}+2\times\frac{1}{4G_{N}}\rho_{b},$ (3.11)
@@
[…]
where in the second step we utilize the transformations in eqs. 3.2, 3.4 and 3.6 to express the entanglement entropy in the AdS black string coordinates. The results agrees with the entanglement entropy in the boundary channel in eq. 2.14, with an additional term arising due to the contributions from the conformal matter on the EOW brane.
Bulk computation : From the bulk perspective the holographic entanglement entropy can be computed using the DES formula eq. 2.9, where the first term can be expressed as the combined lengths of the two RT surfaces $PP^{\prime}$ and $QQ^{\prime}$ denoted by the green curves in fig. 2. The geodesic length between two arbitrary points $Y_{i}$ and $Y_{j}$ in AdS3 may be expressed in the Poincaré coordinates as
$L_{ij}=\cosh^{-1}\left[-\xi_{ij}\right],$ (3.13)
@@
$\xi_{ij}=Y_{i}\cdot Y_{j}=-\frac{(x_{i}-x_{j})^{2}+(\tau_{i}-\tau_{j})^{2}+z_{i}^{2}+z_{j}^{2}}{2z_{i}z_{j}}.$ (3.14)
Using eq. 3.13 the combined length of $PP^{\prime}$ and $QQ^{\prime}$ may be determined as
[…]
Upon extremization of the above expression we obtain $y_{{}_{P^{\prime}}}=x_{1},\tau_{{}_{P^{\prime}}}=\tau_{1}$ and $y_{{}_{Q^{\prime}}}=x_{2},\tau_{{}_{Q^{\prime}}}=\tau_{2}$, which when put back into $S_{gen}$ gives the final holographic entanglement entropy in the island phase (upon utilizing the Brown-Henneaux relation $c=\frac{3}{2G}$) as subsection 3.2.1.
@@ 3.2.2 No-Island Phase
Field theory computation : In this phase the points $P$ and $Q$ are considered to be far away from the brane such that the subsystem $A$ on the radiation bath has no corresponding island region on the brane. This is illustrated in fig. 3. The entanglement entropy in this scenario is once again obtained using eq. 2.8, where the effective entanglement entropy is determined in terms of the 2-point correlator of twist fields at the points $P$ and $Q$ as
$S_{eff}=\lim_{n\to 1}\frac{1}{1-n}\log\langle\sigma_{n}(P)\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}.$ (3.17)
Due to the absence of islands regions on the brane corresponding to the radiation subsystem $A$, the area term in eq. 2.8 does not contribute in this case. As a result of which the generalized entanglement entropy is simply the effective term as $S_{gen}=S_{eff}$. The entanglement entropy is then obtained as
$S(A)=\frac{c}{6}\log\left[\frac{(x_{2}-x_{1})^{2}+(\tau_{2}-\tau_{1})^{2}}{z_{\epsilon_{1}}z_{\epsilon_{2}}}\right],$ (3.18)
where once again using the transformations described in eqs. 3.2, 3.4 and 3.6 may be utilized to express the result in the AdS black string coordinates as
$S(A)=\frac{c}{6}\log\left[\frac{u_{h}}{LR}\left(\Delta_{L}+\Delta_{R}+2\sqrt{\Delta_{L}\Delta_{R}}\cosh\frac{t}{u_{h}}\right)\right]+\frac{c}{3}\rho_{\epsilon}.$ (3.19)
We observe that this result matches exactly with the bulk channel results in eq. 2.14. This is to be expected, since the conformal matter do not contribute to the entanglement entropy due to the absence of any corresponding entanglement entropy island on the brane.
Bulk computation : The holographic entanglement entropy in this case is once again computed using the DES formula given in eq. 2.9, where the first term is expressed in terms of the geodesic length of the Hartman-Maldacena (HM) surface $PQ$ represented by the green curve in fig. 3, which is given as
$S_{RT}=\frac{1}{4G}\cosh^{-1}\left[\frac{(x_{2}-x_{1})^{2}+(\tau_{2}-\tau_{1})^{2}+z_{\epsilon_{1}}^{2}+z_{\epsilon_{2}}^{2}}{2z_{\epsilon_{1}}z_{\epsilon_{2}}}\right].$ (3.20)
As elaborated earlier, due to the absence of corresponding island regions on the EOW brane, second term in eq. 2.9 does not contribute in this case. As a result we have $S_{gen}=S_{RT}$, and the final holographic entanglement entropy is then obtained as eq. 3.19 on application of the Brown-Henneaux relation.
@@ 4.1 Including $T\bar{T}$ deformation
$ds^{2}=\frac{1}{z^{2}}\left(-dt^{2}+dx^{2}+dz^{2}\right),\qquad z>z_{c},\qquad x\geq 0.$ (4.1)
Such a metric is equivalent to considering a tensionless EOW brane at $x=0$, as elaborated in [31, 81]. As the boundary of the $Z_{2}$ quotient $T\bar{T}$ deformed CFT2 moves from $z\to 0$ to some cut-off $z_{c}$, the EOW brane in the bulk now starts from $z=z_{c}$. This modifies the transformations in eq. 3.6 as
[…]
as a result of which the metric in eq. 4.1 becomes
[…]
@@
[…]
where we observe that the conformal factor now differs from that in eq. 2.4 to incorporate the radial cut-off $z_{c}$ (and in turn the $T\bar{T}$ deformation parameter $\mu$ via eq. 2.19).
As elaborated earlier in subsection 2.1, incorporating conformal matter localised on a tensionless EOW brane induces a tension on the brane, which in turn relocates the brane to some constant angle. To impose transparent boundary conditions for the $T\bar{T}$ deformed bath, we include the same field theory on the EOW brane with an identical deformation parameter to that of the $T\bar{T}$ deformed CFT on the cut-off boundary. As discussed in [76, 27] inclusion of this conformal matter to the EOW brane does not backreact on the bulk geometry. This allows us to assume that the cut-off prescription of the holographic dual of the $T\bar{T}$ deformed CFTs is still applicable in the large $N$ approximation.
@@
From eq. 2.19 it follows that the $T\bar{T}$ deformation parameter is related to the finite cut-off in the Poincaré coordinates as
$\mu\propto z_{c}^{2}.$ (4.5)
Using eq. 3.2, this proportionality may be equivalently expressed in terms of the finite cut-off in the AdS black string coordinates as
$\mu\propto e^{-2\rho_{c}}.$ (4.6)

### v2

@@ 2.4 $T\bar{T}$ deformation
We now briefly review $T\bar{T}$ deformation in CFTs and its holographic dual. The $T\bar{T}$ deformation of a $2d$ conformal field theory considered on a flat space is defined as Cavaglia:2016oda ; Smirnov:2016lqw
$\frac{dI_{QFT}^{(\mu)}}{d\mu}=-2\pi\int d^{2}x(T\bar{T})_{\mu},\qquad I_{QFT}^{(\mu)}\Bigg|_{\mu=0}=I_{CFT},$ (15)
@@
$(T\bar{T})=\frac{1}{8}\left(T_{ab}T^{ab}-(T^{a}_{a})^{2}\right).$ (16)
This operator is shown to satisfy the factorization formula as Zamolodchikov:2004ce
$\langle T\bar{T}\rangle=\frac{1}{8}\left(\langle T_{ab}\rangle\langle T^{ab}\rangle-\langle T_{a}^{a}\rangle^{2}\right).$ (17)
The corresponding bulk dual of the $T\bar{T}$ deformed CFT2 in a flat space was proposed by the authors in McGough:2016lol ; Lewkowycz:2019xse to be a quantum gravity in AdS3 with a radial cut-off as
$ds^{2}=\frac{1}{z^{2}}\left(-dt^{2}+dx^{2}+dz^{2}\right),\qquad z>z_{c},$ (18)
where Dirichlet boundary conditions are imposed at $z=z_{c}$. The deformation parameter is related to this radial cut-off $z_{c}$ asLewkowycz:2019xse
$\mu=8G_{N}z_{c}^{2}=\frac{12}{c}z_{c}^{2}.$ (19)
@@ 3.1 Including brane defects
$ds^{2}=d\rho^{2}+\frac{\cosh^{2}\rho}{u^{2}}\bigg[\left(1-\frac{u}{u_{h}}\right)d\tau_{bs}^{2}+\frac{du^{2}}{\left(1-\frac{u}{u_{h}}\right)}\bigg],$ (20)
@@
$\omega=\frac{u_{h}e^{\frac{i\tau_{bs}}{2u_{h}}}}{\sqrt{1-\frac{u}{u_{h}}}},$ (21)
@@
$ds^{2}=d\rho^{2}+\frac{4u_{h}^{2}\cosh^{2}\rho}{(u_{h}^{2}-\omega\bar{\omega})^{2}}d\omega d\bar{\omega}.$ (22)
@@
$\omega=\frac{u_{h}}{\tau_{p}+iy-\frac{i}{2}}-iu_{h},$ (23)
@@
$ds^{2}=d\rho^{2}+\Omega_{y}^{2}(d\tau_{p}^{2}+dy^{2}),\qquad\Omega_{y}=\Big|\frac{\cosh\rho}{y}\Big|,$ (24)
which is the same as eq.˜3. The conformal boundary is now located at $y=0$. Finally using the transformations
$z=y\operatorname{sech}\rho,\qquad x=y\tanh\rho,$ (25)
@@ 3.2 Entanglement Entropy
We now compute the entanglement entropy for a subsystem $A\equiv[L,u_{h}]\cup[u_{h},R]$ on a constant time slice in the radiation baths of the eternal black hole model described previously. The endpoints of $A$ are denoted by the points $P\equiv(t,L,\rho_{\epsilon})$ and $Q\equiv(-t+i2\pi u_{h},R,\rho_{\epsilon})$ on the asymptotic boundary. [^4] Note that to account for the reversal of the timelike Killing vector as one crosses black string horizon from the left to the right partition, we take the time coordinate on the right partition $t_{R}$ to be related to that of the left partition $t_{L}$ as $t_{R}\to-t_{L}+i2\pi u_{h}$. For simplicity of computation we begin in the Poincaré coordinates, where the corresponding endpoints are now at $P\equiv(\tau_{1},x_{1},z_{\epsilon_{1}})$ and $Q\equiv(\tau_{2},x_{2},z_{\epsilon_{2}})$ such that $z_{\epsilon_{1}},z_{\epsilon_{2}}\to 0$. We later transform the final result using eqs.˜21, 23 and 25 to express the entanglement entropy in the AdS black string coordinates once again. Depending on the distance of subsystem $A$ from the EOW brane considered at a constant $\rho=-\rho_{b}$, we come across two distinct entanglement entropy phases, which are discussed below.
@@ 3.2.1 Island Phase
Field theory computation : In this phase the points $P$ and $Q$ are assumed to be close to the brane such that the subsystem $A$ on the radiation bath has a corresponding island region on the brane, as illustrated in fig.˜2. Under these circumstances the entanglement entropy may be determined using the island formula in eq.˜8, where the effective entanglement entropy may be expressed in terms of the 4-point correlators of twist fields at the points $P,P^{\prime},Q$ and $Q^{\prime}$, which in the large $c$ limit factorizes as
$\langle\sigma_{n}(P)\bar{\sigma}_{n}(P^{\prime})\sigma_{n}(Q^{\prime})\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}\approx\langle\sigma_{n}(P)\bar{\sigma}_{n}(P^{\prime})\rangle_{{}_{\text{CFT}}}\langle\sigma_{n}(Q^{\prime})\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}.$ (27)
This allows us to express the first term in eq.˜8 as
$S_{eff}=\lim_{n\to 1}\frac{1}{1-n}\Big[\log\Omega_{P^{\prime}}^{-2h_{n}}\langle\sigma_{n}(P)\bar{\sigma}_{n}(P^{\prime})\rangle_{{}_{\text{CFT}}}+\log\Omega_{Q^{\prime}}^{-2h_{n}}\langle\sigma_{n}(Q^{\prime})\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}\Big],$ (28)
where $\Omega_{P^{\prime}}$ and $\Omega_{Q^{\prime}}$ are the conformal factor at the points $P^{\prime}$ and $Q^{\prime}$ respectively, which are determined using eq.˜24. Assuming $P^{\prime}\equiv(-y_{{}_{P^{\prime}}},\tau_{{}_{P^{\prime}}})$ and $Q^{\prime}\equiv(-y_{{}_{Q^{\prime}}},\tau_{{}_{Q^{\prime}}})$, with the coordinate $y$ running along the brane, the effective entanglement entropy may be obtained using the standard form of the 2-point correlator as
@@
$\qquad\left.+\log\frac{1}{y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b}}+\log\frac{1}{y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b}}\right)$ (29)
where $\epsilon_{y}$ denotes the UV cut-off on the brane. Adding to it the area term provided in Deng:2020ent , the generalized entanglement entropy in this scenario may then be expressed as
$S_{gen}=S_{eff}+2\times\frac{1}{4G_{N}}\rho_{b},$ (30)
@@
$S(A)$ $=\frac{c}{6}\left(\log\frac{2x_{1}}{z_{\epsilon_{1}}}+\log\frac{2x_{2}}{z_{\epsilon_{2}}}+2\rho_{b}+2\log\frac{2}{\epsilon_{y}\operatorname{sech}\rho_{b}}\right)$
$=\frac{c}{3}\left(\rho_{\epsilon}+\rho_{b}+\log\frac{2}{\epsilon_{y}\operatorname{sech}\rho_{b}}\right)$ (31)
where in the second step we utilize the transformations in eqs.˜21, 23 and 25 to express the entanglement entropy in the AdS black string coordinates. The results agrees with the entanglement entropy in the boundary channel in eq.˜14, with an additional term arising due to the contributions from the conformal matter on the EOW brane.
Bulk computation : From the bulk perspective the holographic entanglement entropy can be computed using the DES formula eq.˜9, where the first term can be expressed as the combined lengths of the two RT surfaces $PP^{\prime}$ and $QQ^{\prime}$ denoted by the green curves in fig.˜2. The geodesic length between two arbitrary points $Y_{i}$ and $Y_{j}$ in AdS3 may be expressed in the Poincaré coordinates as
$L_{ij}=\cosh^{-1}\left[-\xi_{ij}\right],$ (32)
@@
$\xi_{ij}=Y_{i}\cdot Y_{j}=-\frac{(x_{i}-x_{j})^{2}+(\tau_{i}-\tau_{j})^{2}+z_{i}^{2}+z_{j}^{2}}{2z_{i}z_{j}}.$ (33)
Using eq.˜32 the combined length of $PP^{\prime}$ and $QQ^{\prime}$ may be determined as
$S_{RT}=$ $\frac{1}{4G}\cosh^{-1}\left[\frac{2x_{1}y_{{}_{P^{\prime}}}\tanh\rho_{b}+x_{1}^{2}+y_{{}_{P^{\prime}}}^{2}+(\tau_{1}-\tau_{{}_{P^{\prime}}})^{2}+z_{\epsilon_{1}}^{2}}{2z_{\epsilon_{1}}y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b}}\right]$
$+\frac{1}{4G}\cosh^{-1}\left[\frac{2x_{2}y_{{}_{Q^{\prime}}}\tanh\rho_{b}+x_{2}^{2}+y_{{}_{Q^{\prime}}}^{2}+(\tau_{2}-\tau_{{}_{Q^{\prime}}})^{2}+z_{\epsilon_{2}}^{2}}{2z_{\epsilon_{2}}y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b}}\right],$ (34)
where using eq.˜25 the coordinates of the bulk points $P^{\prime}$ and $Q^{\prime}$ are expressed as $P^{\prime}\equiv(\tau_{{}_{P^{\prime}}},-y_{{}_{P^{\prime}}}\tanh\rho_{b},y_{{}_{P^{\prime}}}\operatorname{sech}\rho_{b})$ and $Q^{\prime}\equiv(\tau_{{}_{Q^{\prime}}},-y_{{}_{Q^{\prime}}}\tanh\rho_{b},y_{{}_{Q^{\prime}}}\operatorname{sech}\rho_{b})$. The contribution from the defect term in eq.˜9 may be obtained directly from Deng:2020ent , and the generalized entanglement entropy may then be described as
$S_{gen}=S_{RT}+\frac{c}{3}\log\frac{2}{\epsilon_{y}\operatorname{sech}\rho_{b}}.$ (35)
Upon extremization of the above expression we obtain $y_{{}_{P^{\prime}}}=x_{1},\tau_{{}_{P^{\prime}}}=\tau_{1}$ and $y_{{}_{Q^{\prime}}}=x_{2},\tau_{{}_{Q^{\prime}}}=\tau_{2}$, which when put back into $S_{gen}$ gives the final holographic entanglement entropy in the island phase (upon utilizing the Brown-Henneaux relation $c=\frac{3}{2G}$) as section˜3.2.1.
@@ 3.2.2 No-Island Phase
Field theory computation : In this phase the points $P$ and $Q$ are considered to be far away from the brane such that the subsystem $A$ on the radiation bath has no corresponding island region on the brane. This is illustrated in fig.˜3. The entanglement entropy in this scenario is once again obtained using eq.˜8, where the effective entanglement entropy is determined in terms of the 2-point correlator of twist fields at the points $P$ and $Q$ as
$S_{eff}=\lim_{n\to 1}\frac{1}{1-n}\log\langle\sigma_{n}(P)\bar{\sigma}_{n}(Q)\rangle_{{}_{\text{CFT}}}.$ (36)
Due to the absence of islands regions on the brane corresponding to the radiation subsystem $A$, the area term in eq.˜8 does not contribute in this case. As a result of which the generalized entanglement entropy is simply the effective term as $S_{gen}=S_{eff}$. The entanglement entropy is then obtained as
$S(A)=\frac{c}{6}\log\left[\frac{(x_{2}-x_{1})^{2}+(\tau_{2}-\tau_{1})^{2}}{z_{\epsilon_{1}}z_{\epsilon_{2}}}\right],$ (37)
where once again using the transformations described in eqs.˜21, 23 and 25 may be utilized to express the result in the AdS black string coordinates as
$S(A)=\frac{c}{6}\log\left[\frac{u_{h}}{LR}\left(\Delta_{L}+\Delta_{R}+2\sqrt{\Delta_{L}\Delta_{R}}\cosh\frac{t}{u_{h}}\right)\right]+\frac{c}{3}\rho_{\epsilon}.$ (38)
We observe that this result matches exactly with the bulk channel results in eq.˜14. This is to be expected, since the conformal matter do not contribute to the entanglement entropy due to the absence of any corresponding entanglement entropy island on the brane.
Bulk computation : The holographic entanglement entropy in this case is once again computed using the DES formula given in eq.˜9, where the first term is expressed in terms of the geodesic length of the Hartman-Maldacena (HM) surface $PQ$ represented by the green curve in fig.˜3, which is given as
$S_{RT}=\frac{1}{4G}\cosh^{-1}\left[\frac{(x_{2}-x_{1})^{2}+(\tau_{2}-\tau_{1})^{2}+z_{\epsilon_{1}}^{2}+z_{\epsilon_{2}}^{2}}{2z_{\epsilon_{1}}z_{\epsilon_{2}}}\right].$ (39)
As elaborated earlier, due to the absence of corresponding island regions on the EOW brane, second term in eq.˜9 does not contribute in this case. As a result we have $S_{gen}=S_{RT}$, and the final holographic entanglement entropy is then obtained as eq.˜38 on application of the Brown-Henneaux relation.
@@ 4.1 Including $T\bar{T}$ deformation
$ds^{2}=\frac{1}{z^{2}}\left(-dt^{2}+dx^{2}+dz^{2}\right),\qquad\text{such that }z>z_{c},\quad x\geq 0.$ (40)
Such a metric is equivalent to considering a tensionless EOW brane at $x=0$, as elaborated in Kastikainen:2021ybu ; Aharony:2010ay . As the boundary of the $Z_{2}$ quotient $T\bar{T}$ deformed CFT2 moves from $z\to 0$ to some cut-off $z_{c}$, the EOW brane in the bulk now starts from $z=z_{c}$. This modifies the transformations in eq.˜25 as
$z=z_{c}+y\operatorname{sech}\rho,\qquad x=y\tanh\rho,$ (41)
as a result of which the metric in eq.˜40 becomes
$ds^{2}=\frac{1}{\left(z_{c}+y\operatorname{sech}\rho\right)^{2}}\left(d\tau_{p}^{2}+dy^{2}+\frac{y^{2}}{\cosh^{2}\rho}d\rho^{2}\right).$ (42)
@@
$ds_{brane}^{2}=\Omega_{y}^{2}(d\tau_{p}^{2}+dy^{2}),\qquad\Omega_{y}=\Big|\frac{1}{z_{c}+y\operatorname{sech}\rho_{0}}\Big|,$ (43)
where we observe that the conformal factor now differs from that in eq.˜4 to incorporate the radial cut-off $z_{c}$ (and in turn the $T\bar{T}$ deformation parameter $\mu$ via eq.˜19).
As elaborated earlier in section˜2.1, incorporating conformal matter localised on a tensionless EOW brane induces a tension on the brane, which in turn relocates the brane to some constant angle. To impose transparent boundary conditions for the $T\bar{T}$ deformed bath, we include the same field theory on the EOW brane with an identical deformation parameter to that of the $T\bar{T}$ deformed CFT on the cut-off boundary. As discussed in Deng:2023pjs ; Basu:2024xjq inclusion of this conformal matter to the EOW brane does not backreact on the bulk geometry. This allows us to assume that the cut-off prescription of the holographic dual of the $T\bar{T}$ deformed CFTs is still applicable in the large $N$ approximation.
@@
From eq.˜19 it follows that the $T\bar{T}$ deformation parameter is related to the finite cut-off in the Poincaré coordinates as
$\mu\propto z_{c}^{2}.$ (44)
Using eq.˜21, this proportionality may be equivalently expressed in terms of the finite cut-off in the AdS black string coordinates as
$\mu\propto e^{-2\rho_{c}}.$ (45)